# Cliqueformer: Model-Based Optimization with Structured Transformers

### v1

@@ 5 Experiments
In this section, we provide the empirical evaluation of Cliqueformer. We begin by benchmarking Cliqueformer against prior methods on tasks from the MBO literature. Then, we finish by evaluating the benefit of the novel FGM decomposition layer in Cliqueformer through an ablation study.
@@ 5.1 Benchmarking
We compare our model to three classes of algorithms, each represented by a proven prior method. As a naïve baseline, we employ gradient ascent on a learned model (Grad. Asc.). To compare to exploratory methods, we use Reward-Weighted Regression (peters2007reinforcement, RWR) which learns by regressing the policy against its most promising perturbations. To represent the recently proposed conservative algorithms, we compare to state-of-the-art Conservative Objective Models (trabucco2021conservative; kumar2021data, COMs). Additionally, to evaluate the importance of the novel elements of our architecture and training, we evaluate gradient ascent with the transformer backbone (vaswani2017attention, Transformer). For COMs and other methods, we use the recommended hyperparameter setting from (trabucco2021conservative). While Cliqueformer can be tuned for each task, we keep most of the hyperparameters the same: refer to Appendix B for more details.
For every method, we report an empirical estimate of its $100^{th}$ percentile, similarly to trabucco2021conservative. We estimate it by averaging the values of the top 10 designs out of 1000 candidates, averaged across 5 seeds. In the following paragraphs, we introduce benchmark tasks, from MBO literature, that we use in our experiments, and analyze their results.
@@
This is a suite of tasks designed to expose vulnerability of MBO models (grudzien2024functional). The data pairs $({\mathbf{x}},{\textnormal{y}})$ are generated by first drawing a standard normal vector ${\mathbf{z}}\sim N(0_{d_{z}},I_{d_{z}})$, where $d=\{11,31,41,61\}$, then computing y as a sum of radial-basis functions of $d_{C}$-dimensional cliques of a pre-defined FGM. The observed vector ${\mathbf{x}}$ is a non-linear transformation of ${\mathbf{z}}$, i.e., ${\mathbf{x}}=T({\mathbf{z}})\in\mathcal{T}\subset\mathbb{R}^{d}$, where $d>d_{z}$, while ${\mathbf{z}}$ itself is hidden from the data. Such tasks allow us to study whether an MBO method learns to produce valid designs by verifying that ground-truth inputs ${\mathbf{z}}$ can be recovered from them. That is, a design $\hat{{\mathbf{x}}}$ for which the map $T^{-1}(\hat{{\mathbf{x}}})$ is ill-defined is considered invalid. In our experiments, an invalid design receives a value of $\infty$- ∞. We report the average validity of designs produced by the methods in blue. The results in Table 1 show that the most able method at keeping its proposals at the manifold of valid designs is Cliqueformer. Importantly, this ability does not diminish even in higher-dimensional tasks, while the second-best such method, COMs, gradually loses this ability. We also use these tasks to study if a model is capable of exploiting the RBF’s structure in the optimization step by varying the effective dimensionality $d_{z}$ of the data while keeping $d_{C}$ fixed. While the naive, the exploratory, and Transformer baselines perform poorly on these tasks overall, COMs’s performance clearly drops as the task dimension increases. Meanwhile, as predicted by Theorem 1, Cliqueformer attains similar, strong performance across all tasks.
[…]
Lat. RBF 11 $\infty$- ∞ $\pm$ 0.00 (0%) 0.08 $\pm$ 0.08 (1%) 0.66 $\pm$ 0.04 (72%) 0.47 $\pm$ 0.05 (60%) 0.65 $\pm$ 0.07 (74%)
Lat. RBF 31 $\infty$- ∞ $\pm$ 0.00 (0%) 0.31 $\pm$ 0.10 (3%) 0.50 $\pm$ 0.05 (32%) $\infty$- ∞ $\pm$ 0.00 (0%) 0.64 $\pm$ 0.05 (76%)
Lat. RBF 41 $\infty$- ∞ $\pm$ 0.00 (0%) 0.35 $\pm$ 0.08 (3%) 0.45 $\pm$ 0.06 (16%) 0.20 $\pm$ 0.01 (75%) 0.66 $\pm$ 0.05 (72%)
Lat. RBF 61 $\infty$- ∞ $\pm$ 0.00 (0%) 0.29 $\pm$ 0.10 (4%) 0.25 $\pm$ 0.04 (7%) 0.16 $\pm$ 0.03 (64%) 0.66 $\pm$ 0.05 (68%)
Superconductor 1.13 $\pm$ 0.08 1.03 $\pm$ 0.07 0.97 $\pm$ 0.08 0.96 $\pm$ 0.05 1.43 $\pm$ 0.04
TF-Bind-8 0.99 $\pm$ 0.00 1.58 $\pm$ 0.03 1.57 $\pm$ 0.02 1.48 $\pm$ 0.03 1.58 $\pm$ 0.01
DNA HEPG2 2.16 $\pm$ 0.07 1.91 $\pm$ 0.12 1.20 $\pm$ 0.09 2.13 $\pm$ 0.06 2.10 $\pm$ 0.07
DNA k562 2.11 $\pm$ 0.06 1.91 $\pm$ 0.11 1.80 $\pm$ 0.12 2.60 $\pm$ 0.19 3.15 $\pm$ 0.07
Ave.score($\uparrow$) 0.80 0.93 0.93 0.80 1.46
Ave.rank($\downarrow$) 3.75 3.00 3.13 3.50 1.38
Table 1: Experimental results of Cliqueformer and the baselines. Each score is the mean of values of $TopK$=10 of $B$=1000 designs, averaged over 5 runs. The values were normalized with the min-max scheme, where the minimum and the maximum are taken from the dataset. We also provide the standard deviation estimates as the standard deviation for each of the $TopK$ samples, averaged over the runs. Additionally, for Lat. RBF tasks, we provide the average validity in blue, calculated as the percentage of valid designs from the $B$ produced candidates, averaged across the runs. We provide average score (the higher the better) and the average rank for each method. For the average score $\infty$- ∞ was taken into calculation as zero.
@@
This task poses a challenge of designing a superconducting material, represented by an 81-dimensional vector, with as high critical temperature as possible (hamidieh2018data). It tests abilities of MBO models in real-world continuous problems. As the very high score of gradient ascent shows, in contrast to Lat. RBF, this task rewards greedy optimizer updates. While we developed Cliqueformer paying attention to distribution shift, its ability to compose individual in-distribution cliques into a joint solution (stitching) allows for large improvements that outperform all baselines. In particular, it greatly outperforms COMs whose conservative objective caps the extent of improvement in the design scores, and Transformer which does not attain Cliqueformer’s decomposition.
@@
In these discrete tasks we optimize DNA sequences of length 8 and 200, respectively. In TFBind-8 the target is the sequence’s binding affinity with a particular transcription factor, while in DNA Enhancers we maximize HEPG2 and k562 activity levels. Being low-dimensional, TFBind-8 (trabucco2022design) is a testbed that allows us to verify MBO models’ ability to solve discrete tasks. Cliqueformer solves this problem very efficiently, largely improving upon the dataset, but it is worth noting that most baselines, with the exception of gradient ascent with and without Transformer backbone, performed similarly. Thus, while the TFBind-8 task does not favor greedy design optimization, Cliqueformer is still able to leverage its other strengths to achieve great performance. Then, we use the DNA Enhancers tasks (uehara2024understanding) to study the scalability of our method to very high-dimensional problems and large datasets (approximately $2\times 10^{5}$ examples in this case). The results in Table 1 and the high score of gradient ascent show that these tasks favor direct, greedy optimization of designs more than they benefit from conservatism of COMs. Nevertheless, Cliqueformer is able to greatly exceed the quality of observed designs, and performs on par with gradient ascent in HEPG2, and greatly outperforms all baselines in k562, confirming its ability to learn structure within discrete and very high-dimensional data. Overall, averaging across all tested tasks, Cliqueformer achieves the best overall performance.
5.2 Ablation
@@
While the decomposing mechanism of Cliqueformer is entirely novel, other components, such as transformer blocks (vaswani2017attention) and variational-information bottlenecks (kingma2013auto; alemi2016deep) are known and powerful deep learning tools. In this subsection, we verify the utility of the decomposing component with an ablation study, in which we sweep over the number of cliques of Cliqueformer for a few representative tasks. In each of the tested tasks we fix the size of the latent variable ${\mathbf{z}}$ and sweep over the number of cliques $N_{\text{clique}}$ into which it can be decomposed. We cover the case $N_{\text{clique}}=1$ to compare Cliqueformer to an FGM-oblivious VAE with transformer backbone and AdamW design optimizer.
(a) TFBind-8
(b) Superconductor
(c) Lat. RBF 41
Figure 5: Ablation experiments on the number of cliques used in the FGM decomposition of Cliqueformer. For each task, we fix the size of the latent variable and the overlap size, and sweep over possible $N_{\text{clique}}$ values. We use TFBind-8 (left), Superconductor (center), and Lat. RBF 41 (right) tasks. The x-ais (log-scale) denotes the number of cliques, and the y-axis is the final score of the model.
@@
Results in Figure 5 show that Cliqueformer does benefit from the FGM decomposition in all tested tasks. However, the results suggest that the optimal number of cliques varies between tasks, e.g. 4 for TFBind-8 and Superconductor, but 2 for Lat. RBF 41. In our experiments, we consistently obtained good performance by setting the clique size to $d_{\text{clique}}=3$ and choosing the number of cliques so that the total latent dimension approximately matches that of the design. In DNA Enhancers, we doubled the clique size and halved the number of cliques to decrease the computational cost of attention. More details of hyperparameters can be found in Appendix B. An in-depth analysis of the relation between hyperparameters and the performance is an exciting avenue of future work.
@@ 6 Conclusion
In this work, we proposed Cliqueformer—a scalable neural model for model-based optimization. We derived its building blocks following recent advances in MBO theory centered in functional graphical models, equipping it with an ability to acquire structure of the black-box target function. This ability sets the model free from requiring explicit conservative regularization or iterative retraining to propose in-distribution designs. Empirically, Cliqueformer achieves state-of-the-art performance, which it sustains across all tested tasks. Cliqueformer opens an exciting avenue of research in MBO focused on scaling design datasets and large neural networks.
@@ Acknowledgements
Appendix
Appendix A Theoretical details
@@
where $\delta$ is the PAC error probability (shalev2014understanding).
@@
Since the density of ${\mathbf{x}}$ is positive and continuous, we can form a bijection that maps ${\mathbf{x}}$ to another random variable ${\mathbf{z}}\in\mathbb{R}^{l}$, where $l\leq d$, that follows the standard-normal distribution (dai2019diagnosing, Appendix E). We denote this bijection as $Z({\mathbf{x}})$. Let us define
@@
is also a function of all others (grudzien2024functional, Lemma 1). Consider now a rotation $\rho:{\mathbf{z}}\mapsto{\mathbf{v}}=({\textnormal{v}}_{1},\dots,{\textnormal{v}}_{l})$ such that ${\textnormal{v}}_{1}=\frac{1}{\sqrt{l}}\sum_{i=1}^{l}{\textnormal{z}}_{i}$. Then, ${\mathbf{v}}\sim N(0_{l},I_{l})$, and y can be expressed in terms of ${\mathbf{v}}$ as ${\textnormal{y}}=f^{v}({\mathbf{v}})=\exp({\textnormal{v}}_{1})$. Then, the FGM of $f^{v}$ has no edges, since it depends on only one variable, inducing no interactions between any two variables. Recall that ${\mathbf{x}}=Z^{-1}({\mathbf{z}})$. Then, ${\mathbf{x}}$ be represented by standard-normal ${\mathbf{z}}$ and ${\mathbf{v}}$, obtainable by
@@
Appendix B Experimental details

### v2

@@ 5 Experiments
In this section, we provide the empirical evaluation of Cliqueformer. We begin by benchmarking Cliqueformer against prior methods on tasks from the MBO literature. We then evaluate the effect of the novel FGM decomposition layer in Cliqueformer through an ablation study.
@@ 5.1 Benchmarking
We compare our model to five classes of algorithms, each represented by a proven prior method. As a naïve baseline, we employ gradient ascent on a learned model (Grad. Asc.). To compare to exploratory methods, we use Reward-Weighted Regression (Peters & Schaal, 2007, RWR) which learns by regressing the policy against its most promising perturbations. To represent conservative algorithms, we compare to state-of-the-art Conservative Objective Models (Trabucco et al., 2021; Kumar et al., 2021, COMs). As a conditional generative modeling in MBO baseline, we use Denoising Diffusion Optimization Models (Krishnamoorthy et al., 2023, DDOM) which is based on diffusion models (Ho et al., 2020). Additionally, to evaluate the importance of the novel elements of our architecture and training, we evaluate gradient ascent with the transformer backbone (Vaswani et al., 2017, Transformer). We use standard hyperparameters for the baselines (Trabucco et al., 2021; Krishnamoorthy et al., 2023). While Cliqueformer can be tuned for each task, we keep most of the hyperparameters the same: see Appendix D for more details.
Following prior work (Trabucco et al., 2022; Krishnamoorthy et al., 2023), we normalize scores to $[0,1]$ based on the dataset’s range and report the empirical top 1%, estimated by averaging the top 10 designs from 1000 candidates across 5 seeds. Below, we present our benchmark tasks and analyze the performance of various approaches.
(a) TFBind-8 ($d_{z}$ fixed)
(b) Superconductor ($d_{z}$ fixed)
(c) TFBind-8 ($d_{\text{clique}}$ fixed)
(d) Superconductor ($d_{\text{clique}}$ fixed)
Figure 6: Ablation experiments on the number of cliques $N_{\text{clique}}$ used in the FGM decomposition of Cliqueformer (the $y$-axis shows the design score and the $x$-axis shows the number of cliques used). In Figures (6(a)) & (6(b)), for each task, we fixed the size $d_{z}$ of the latent variable ${\mathbf{z}}$ and varied the number of cliques $N_{\text{clique}}$. In Figures (6(c)) & (6(d)), for each task, we fixed the clique size $d_{\text{clique}}$, and varied the number of cliques $N_{\text{clique}}$. In every figure, the red bar corresponds to the best (maximal) score.
@@
This suite of tasks is designed to stress-test MBO methods (Grudzien et al., 2024). The data pairs $({\mathbf{x}},{\textnormal{y}})$ are generated by first drawing a standard normal vector ${\mathbf{z}}\sim N(0_{d_{z}},I_{d_{z}})$, where $d_{z}=\{11,31,41,61\}$, then computing y as a sum of radial-basis functions of $d_{C}$-dimensional cliques of a pre-defined FGM. The observed vector ${\mathbf{x}}$ is a non-linear transformation of ${\mathbf{z}}$, i.e., ${\mathbf{x}}=T({\mathbf{z}})\in\mathcal{T}\subset\mathbb{R}^{d}$, where $d>d_{z}$, while ${\mathbf{z}}$ itself is hidden from the data. Such tasks allow us to study whether an MBO method learns to produce valid designs by verifying that ground-truth inputs ${\mathbf{z}}$ can be recovered from them. That is, a design $\hat{{\mathbf{x}}}$ for which the map $T^{-1}(\hat{{\mathbf{x}}})$ is ill-defined is considered invalid. In our experiments, an invalid design receives a value of 00. We report the average validity of designs produced by the methods in parentheses. The results in Table 1 show that the most able method at keeping its proposals at the manifold of valid designs is Cliqueformer. Importantly, this ability does not diminish even in higher-dimensional tasks, while the second-best such method, COMs, gradually loses this ability. We also use these tasks to study if a model is capable of exploiting the RBF’s structure in the optimization step by varying the effective dimensionality $d_{z}$ of the data while keeping $d_{C}$ fixed. While other baselines perform poorly on these tasks overall, COMs’s performance is initially impressive but clearly drops as the task dimension increases. Meanwhile, as predicted by Theorem 1, Cliqueformer sustains strong performance across all tasks.
@@
This task poses a challenge of designing a superconducting material, represented by an 81-dimensional vector, with as high critical temperature as possible (Hamidieh, 2018; Fannjiang & Listgarten, 2020; Trabucco et al., 2022). It tests abilities of MBO models in real-world continuous problems. As the very high score of gradient ascent shows, in contrast to Lat. RBF, this task rewards greedy optimizer updates. While we developed Cliqueformer paying attention to distribution shift, its ability to compose individual in-distribution cliques into a joint solution (stitching) allows for large improvements that outperform all baselines. In particular, it greatly outperforms COMs whose conservative objective caps the extent of improvement in the design scores, and Transformer which does not attain Cliqueformer’s decomposition.
@@
In these discrete tasks we optimize DNA sequences of length 8 and 200, respectively (Trabucco et al., 2022; Uehara et al., 2024b). In TFBind-8, the target is the sequence’s binding affinity with a particular transcription factor, while in DNA Enhancers we maximize HEPG2 and k562 activity levels. The low-dimensional TFBind-8 tasks allows us to evaluate whether a particular method is suitable for solving even simple discrete tasks. Cliqueformer solves this problem very efficiently, largely improving upon the dataset, but it is worth noting that most baselines, with the exception of gradient ascent with and without Transformer backbone, performed similarly. Thus, while the TFBind-8 task does not favor greedy design optimization, Cliqueformer is still able to leverage its other strengths to achieve great performance. Then, we use the DNA Enhancers tasks (Uehara et al., 2024a) to study the scalability of our method to high-dimensional problems and large datasets (approximately $2\times 10^{5}$ examples in this case). The results in Table 1 and the high score of gradient ascent show that these tasks favor direct, greedy optimization of designs more than they benefit from the conservative strategy used by COMs. Nevertheless, Cliqueformer is able to greatly exceed the quality of observed designs, and performs on par with gradient ascent in HEPG2, and greatly outperforms all baselines in k562, confirming its ability to solve high-dimensional MBO problems. Overall, averaging across all tested tasks, Cliqueformer achieves the best overall performance.
5.2 Ablations
@@
While some components of Cliqueformer, such as transformer blocks (Vaswani et al., 2017) and variational-information bottlenecks (Kingma & Welling, 2013; Alemi et al., 2016), are well-known deep learning tools, the decomposing mechanism of our model is novel. In Figures (6(a)) & 6(b)), we verify the utility of the decomposing mechanism with an ablation study, in which we sweep over the number of cliques $N_{\text{clique}}$ of Cliqueformer in the discrete TFBind-8 and continuous Superconductor tasks. In each of the tasks, we fix the size $d_{z}$ of the latent variable ${\mathbf{z}}$ so that it approximately matches the dimensionality of the data, and sweep over the number of cliques into which it can be decomposed under the formula $d_{z}=d_{\text{knot}}+N_{\text{clique}}(d_{\text{clique}}-d_{\text{knot}})$. We cover the case $N_{\text{clique}}=1$ to compare Cliqueformer to an FGM-oblivious VAE with transformer backbone.
Results in Figures (6(a)) & (6(b)) demonstrate the essential role of FGM decomposition in achieving superior performance across both tasks. The results reveal a clear trade-off in parameter selection: with fixed $d_{z}$, too few cliques (particularly $N_{\text{clique}}=1$) prevent effective function decomposition, while too many cliques oversimplify the target function. We further investigated the impact of varying the number of cliques while maintaining fixed clique sizes ($d_{\text{clique}}=3$ for TFBind-8 and $d_{\text{clique}}=21$ for Superconductor). Figures (6(c)) & (6(d)) confirm that the optimal $N_{\text{clique}}$ values remain unchanged, likely because deviating from these values results in latent spaces that are either too constrained or too expansive relative to the data dimensionality.
@@
In our experiments, we consistently obtained good performance by setting the clique size to $d_{\text{clique}}=3$ and choosing the number of cliques so that the total latent dimension approximately matches that of the design. In DNA Enhancers, we doubled the clique size and halved the number of cliques to decrease the computational cost of attention layers. We offer more ablation studies in Appendix C, and we include a detailed report of hyperparameters in Appendix D. An in-depth analysis of the relation between hyperparameters and the performance is an exciting avenue of future work.
@@ 6 Conclusion
In this work, we introduced Cliqueformer, a scalable neural architecture for offline model-based optimization that leverages functional graphical models to learn the structure of black-box target functions. By incorporating recent theoretical advances in MBO, our model eliminates the need for explicit conservative regularization or iterative retraining when proposing designs. Cliqueformer achieves and maintains state-of-the-art performance across all evaluation tasks, paving the way for research into scaling MBO with large neural networks and expanded design datasets.
Impact Statement
Cliqueformer demonstrates how the strengths of generative models, such as transformers, can be leveraged in the new and growing field of offline model-based optimization (MBO). We hope to inspire and lay the foundation for future works in MBO that will effectively capitalize on this link between these two paradigms. We expect such methods, including Cliqueformer, to have a transformative impact on fields where innovation is traditionally driven by extensive wet-lab experiments, such as material science and protein engineering. At the same time, we must remain vigilant as a community, ensuring that these powerful models are developed and deployed responsibly, mitigating potential misuse, such as the creation of harmful substances. By fostering collaboration and ethical practices, we can harness the potential of these tools to accelerate progress and drive meaningful, positive change in society.
@@ Acknowledgements
Appendix A More Related Work
Several reinforcement learning (RL) approaches have been explored extensively for biological sequence design. DyNA-PPO (Angermueller et al., 2019) leverages proximal policy optimization (Schulman et al., 2017) with a model-based variant to improve sample efficiency in the low-round setting typical of wet lab experiments. PEX, also resembling the PPO (Schulman et al., 2017) learning style, (Ren et al., 2022) prioritizes local search through directed evolution (Arnold, 1998) while using a specialized architecture for modeling fitness landscapes. FBGAN (Gupta & Zou, 2018) introduces a feedback loop mechanism to optimize synthetic gene sequences using an external analyzer. However, these methods fundamentally rely on active learning and iterative refinement through oracle queries - DyNA-PPO requires simulator fitting on new measurements, PEX conducts proximal exploration, and FBGAN uses feedback loops with an external analyzer. This makes them unsuitable for offline MBO settings where no additional queries are allowed. Furthermore, while these approaches are specialized for biological sequences, offline MBO aims to tackle a broader class of design problems through static dataset learning.
Appendix B Theoretical details
@@
where $\delta$ is the PAC error probability (Shalev-Shwartz & Ben-David, 2014).
@@
Since the density of ${\mathbf{x}}$ is positive and continuous, we can form a bijection that maps ${\mathbf{x}}$ to another random variable ${\mathbf{z}}\in\mathbb{R}^{l}$, where $l\leq d$, that follows the standard-normal distribution (Dai & Wipf, 2019, Appendix E). We denote this bijection as $Z({\mathbf{x}})$. Let us define
@@
is also a function of all others (Grudzien et al., 2024, Lemma 1). Consider now a rotation $\rho:{\mathbf{z}}\mapsto{\mathbf{v}}=({\textnormal{v}}_{1},\dots,{\textnormal{v}}_{l})$ such that ${\textnormal{v}}_{1}=\frac{1}{\sqrt{l}}\sum_{i=1}^{l}{\textnormal{z}}_{i}$. Then, ${\mathbf{v}}\sim N(0_{l},I_{l})$, and y can be expressed in terms of ${\mathbf{v}}$ as ${\textnormal{y}}=f^{v}({\mathbf{v}})=\exp({\textnormal{v}}_{1})$. Then, the FGM of $f^{v}$ has no edges, since it depends on only one variable, inducing no interactions between any two variables. Recall that ${\mathbf{x}}=Z^{-1}({\mathbf{z}})$. Then, ${\mathbf{x}}$ be represented by standard-normal ${\mathbf{z}}$ and ${\mathbf{v}}$, obtainable by
@@
Computational complexity.
Below (Table 2), we list the computational complexities, as the order of the number of FLOPs, for each method’s training step and design optimization phase, as a function of batch size $B$, number of model layers $L$, model’s hidden dimension $H$, number of exploratory perturbations $P$, number of adversarial training sub-steps $A$, number of design optimization steps $T$, and the number of cliques in an FGM-based model $C$. Note that the majority of quadratic terms, such as $H^{2}$ and $C^{2}$ do not influence runtime much if parallelized on a GPU/TPU. We print in bold terms, such as $\mathcal{\mathbb{T}}$, that contribute to the complexity with sequential operations, thus inevitably affecting the runtime.
Method Training Step Complexity Design Complexity
Grad Asc. $\mathcal{O}(BLH^{2})$ $\mathcal{O}(\mathbf{T}LH^{2})$
RWR $\mathcal{O}(BLH^{2})$ $\mathcal{O}(\mathbf{T}PLH^{2})$
COMs $\mathcal{O}((\mathbf{A}+B)LH^{2})$ $\mathcal{O}(\mathbf{T}LH^{2})$
DDOM $\mathcal{O}(BLH^{2})$ $\mathcal{O}(\mathbf{T}LH^{2})$
Transformer $\mathcal{O}(BLHD(H+D))$ $\mathcal{O}(\mathbf{T}LHD(H+D))$
Cliqueformer $\mathcal{O}(BLH(D(H+D)+C(H+C)))$ $\mathcal{O}(\mathbf{T}LH(D(H+D)+C(H+C)))$
Table 2: Computational complexities (in terms of FLOPs) of methods from Section 5.
Appendix C Extra Experiments
Ablations.
We ablate the components of our model, and its design optimization algorithm, that served as regularizers and were inspired theoretically. That is, we examine the VIB term in the loss from Equation (3.2), and the weight decay step in Algorithm 1. We have not run the ablations for DNA Enhancers tasks due to their high computational cost - around 11 hours on a Google TPU v3-8 for a single run. The results are reported in Table 3.
[…]
Base 0.65 $\pm$ 0.07 0.64 $\pm$ 0.05 0.66 $\pm$ 0.05 0.66 $\pm$ 0.05 1.43 $\pm$ 0.04 1.58 $\pm$ 0.01 0.94 1.83
No VIB 0.64 $\pm$ 0.06 0.68 $\pm$ 0.05 0.67 $\pm$ 0.04 0.66 $\pm$ 0.05 0.92 $\pm$ 0.02 1.47 $\pm$ 0.06 0.84 2.17
No Weight Decay 0.66 $\pm$ 0.06 0.69 $\pm$ 0.06 0.64 $\pm$ 0.07 0.63 $\pm$ 0.05 1.45 $\pm$ 0.04 1.48 $\pm$ 0.06 0.93 2.00
Table 3: Comparison of Cliqueformer (the Base version) to ablative versions—one without the VIB term (No VIB) in Equation 3.2, and one where the model was trained just like the base version but the design optimizer does not use weight decay (No Weight Decay).
We observe that taking out the VIB term can significantly harm the performance, as showcased by the Superconductor task. If the term is kept but the weight decay step is removed from the design optimizer, the final results slightly underperform the full model and algorithm. Thus, we recommend to keep these regularizers.
Different metric.
In the main body of the paper, the results in Table 1 are our estimates of the expected top 1% of the distribution of the design values, computed by taking the average of top 10 values from a batch of 1000 designs, averaged over 5 random seeds. Our rationale behind it was to establish a more concrete metric of evaluation, like the expected top 1%, instead of reporting the maximum of 128 designs that is more common in recent literature (Trabucco et al., 2022; Krishnamoorthy et al., 2023). In fact, we argue that top 1 of 128 is an estimate of the top 1% too (since 1/128=0.78%), but one with more bias and variance. However, for completeness, we report results with this metric (top 1 of 128, averaged over random seeds) in Table 4.
Task Grad.Asc. RWR COMs DDOM Transformer Cliqueformer
Lat. RBF 11 0.00 0.00 0.74 0.00 0.48 0.72
Lat. RBF 31 0.00 0.00 0.49 0.00 0.00 0.65
Lat. RBF 41 0.00 0.00 0.41 0.00 0.00 0.66
Lat. RBF 61 0.00 0.00 0.00 0.00 0.00 0.62
Superconductor 1.19 1.00 0.92 1.20 0.91 1.42
TF-Bind-8 0.99 1.60 1.56 1.55 1.52 1.48
Ave.score $\uparrow$ 0.36 0.43 0.69 0.46 0.49 0.93
Ave.rank $\downarrow$ 3.50 2.83 2.33 2.83 3.50 1.82
Table 4: Experimental results of Cliqueformer and the baselines. Each score is the maximum of values of batch_size=128 designs, averaged over 5 runs. The values were normalized with the min-max scheme, where the minimum and the maximum are taken from the dataset, so that the scores of the designs in the dataset are in range $[0,1]$. We note that, unlike Trabucco et al. (2022), we take the maximum from the data available for the MBO model training (the union of the train and test data), and not from the oracle training data, to make the results more interpretable (see Appendix D). We also provide the standard deviation estimates as the standard deviation for each of the top_k samples, averaged over the runs. We provide average score (the higher the better) and the average rank (the lower the better) for each method.
We observe very similar scores to those in Table 1, but with the variance effect being visible: for example, RWR get poorer results in Lat. RBF 11 (0.00 as compared to 0.08), but stronger in TFBind8 (1.60 as compared to 1.58).
Appendix D Experimental details